Problem Solving: Show your complete Solution.

  1. In an atom an electron moves about its nucleus in a circular path of radius 5.30 x 10-11 m, with a speed of 2.20 x 106 m/s. Find the acceleration of the electron and the centripetal force. (mass of electron = 9.11 x 10-31 kg)


2.A ball weighing 1.5 kg is whirled in a circular path at a speed of 12.0 m/s. If the radius of the circle is 5.6 m, what is the centripetal force?



1
Expert's answer
2021-12-09T09:14:59-0500

1) "a=v^2/R=(2.2\cdot10^6)^2/(5.3\cdot10^{-11})=9.13\cdot10^{22} \ (m/s^2)"


"F=ma=9.13\cdot10^{22}\cdot9.11\cdot10^{-31}=83.17\cdot10^{-9}\ (N)"


2) "F=mv^2/R=1.5\cdot12^2/5.6=38.57\ (N)"